# SpeechCap: Leveraging Playful Impact Captions to Facilitate Interpersonal Communication in Social Virtual Reality

### v1

@@ Abstract.
Social Virtual Reality (VR) emerges as a promising platform bringing immersive, interactive, and engaging mechanisms for collaborative activities in virtual spaces. However, interpersonal communication in social VR is still limited with existing mediums and channels. To bridge the gap, we propose a novel method for mediating real-time conversation in social VR, which uses impact captions, a type of typographic visual effect widely used in videos, to convey both verbal and non-verbal information. We first investigated the design space of impact captions by content analysis and a co-design session with four experts. Next, we implemented SpeechCap as a proof-of-concept system, with which users can communicate with each other using speech-driven impact captions in VR. Through a user study (n=14), we evaluated the effectiveness of the visual and interaction design of impact captions, highlighting the interactivity and the integration of verbal and non-verbal information in communication mediums. Finally, we discussed topics of visual rhetoric, interactivity, and ambiguity as the main findings from the study, and further provided design implications for future work for facilitating interpersonal communication in social VR.
@@
copyright: ccccs: Human-centered computing Virtual realityccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in interaction design
Figure 1. SpeechCap showcases a novel way to mediate voice conversations in social VR by converting speech into playful impact captions. (A) From real-time speech, SpeechCap generates impact captions with customized visual design that conveys both verbal and non-verbal information. The caption “happy” is in a warm and bright color for the positive emotions, and a “smiling” emoji further highlight the pleasant and relaxing atmosphere. (B) Users can play with impact captions like real-world objects, e.g., throwing a “cake”. (C) Once the flying “cake” crashes on another user, it will explode like a firework.
@@ 1. Introduction
With the increasing availability of consumer-level virtual reality head-mounted displays (HMDs)  (Anthes et al., 2016; Li et al., 2021), Social Virtual Reality (VR) emerges as a promising platform that brings immersive, interactive, and engaging experiences to users for a wide range of collaborative activities. Social VR now supports remote conferencing  (Abdullah et al., 2021), immersive learning  (Thanyadit et al., 2022; Peng et al., 2022), healthcare  (Li et al., 2020; Udapola, 2022), collaborative design  (Mei et al., 2021), music composing  (Men and Bryan-Kinns, 2022) and daily activities such as sleeping  (Maloney and Freeman, 2020) and even drinking  (Chen et al., 2024). Among these diverse application areas, social VR helps users overcome physical dispersion, enabling them to communicate and collaborate with each other in the same virtual environment  (Li et al., 2021; Palmer, 1995) by offering channels supporting both verbal and non-verbal communication  (Dzardanova et al., 2022; McVeigh-Schultz and Isbister, 2021; Wei et al., 2024). Text messaging and voice chat are two channels commonly used for verbal communication. Pictures  (Li et al., 2019), avatars  (Fu et al., 2023; Baker et al., 2021; Kolesnichenko et al., 2019; Freeman and Maloney, 2021), and embodied interactions  (Maloney et al., 2020; Wieland et al., 2022) are channels that can convey non-verbal information in social VR.
Yet, these existing communication channels in social VR are preliminary, as they are either borrowed from traditional communication media (e.g., texts, voices, and pictures) lacking adaptations to the nature of social VR  (Montoya Esquer and Lara López, 2023; Rzayev et al., 2021; Dzardanova et al., 2022; Hsieh et al., 2020), or difficult to be reproduced very well by current VR technologies (e.g., gestures, facial expressions, embodied actions)  (Sykownik et al., 2023; Tanenbaum et al., 2020; Wei et al., 2024). Specifically, texts as a communication medium in VR still suffer from difficulties in user inputting  (Montoya Esquer and Lara López, 2023) and readability issues in presentations  (Rzayev et al., 2021; Hsieh et al., 2020). Voice-based conversation in VR can easily be disturbed by sound intrusions from the environment  (Akselrad et al., 2023). Furthermore, non-verbal communication mediums, such as gestures and facial expressions, cannot be well-simulated with the avatars in current VR applications  (Baker et al., 2021; Aseeri and Interrante, 2021; Freeman and Maloney, 2021), and users may find it hard to achieve specific actions or movements using controllers with joysticks and buttons  (Sykownik et al., 2023; Tanenbaum et al., 2020; Li et al., 2019).
@@
To implement and evaluate our impact-caption-inspired approach, we first explored the design space of impact captions regarding the visual appearances and interactivity through a content analysis of a collection of TV show videos. Based on the design space, we held a co-design session with 4 experts (i.e., HCI researchers) to further derive specific design goals for building a system to demonstrate our idea. Following the findings so far, we developed SpeechCap, a proof-of-concept system that takes real-time speech as input to generate interactive impact captions with customized visual appearances in a multi-user VR environment. Using SpeechCap, we conducted an in-lab study with 14 participants to assess impact captions as a medium for supporting interpersonal communication in social VR.
@@
In summary, this research contributes: (i) A comprehensive design space of impact captions in terms of visual and interaction design for supporting interpersonal communication; (ii) A proof-of-concept system, SpeechCap, that enables users to conduct real-time conversations in social VR with the help of interactive impact captions; (iii) An evaluation study of SpeechCap that demonstrates the effectiveness of our impact-caption-inspired approach for communication, highlighting the value of employing creative mediums for supporting interpersonal communication in social VR.
@@ 2.1. Interpersonal Communication in Social Virtual Reality
Social Virtual Reality (VR) refers to immersive virtual spaces where multiple users can synchronously interact with each other through VR head-mounted displays (HMDs)  (McVeigh-Schultz et al., 2019; Freeman and Maloney, 2021). It provides high-fidelity physical presence to facilitate various forms of interpersonal communication with both verbal and non-verbal cues over other computer-mediated communication channels  (Yassien et al., 2020; McVeigh-Schultz and Isbister, 2021). With an immersive and synchronous manner, social VR can effectively support collaborative tasks and promote social interactions in a wide range of application scenarios, including collaborative prototyping  (Mei et al., 2021), healthcare and treatment  (Li et al., 2020; Udapola, 2022), immersive learning  (Thanyadit et al., 2022; Peng et al., 2022; Jensen and Konradsen, 2018), intimate relationship building  (Wang et al., 2023; Freeman and Acena, 2021), and inter-generational communication  (Shen et al., 2024; Wei et al., 2023; Baker et al., 2019).
Although social VR has been increasingly applied in various domains, it remains far from being a socially and emotionally fulfilling digital space for users’ socialization needs  (Wei et al., 2024; Tanenbaum et al., 2020; Maloney et al., 2020). Previous research identified several limitations regarding interpersonal communication in social VR. A known issue is that current social VR applications do not allow users to easily express emotions (i.e., moods and excitements) and intentions through non-verbal cues that are common in real-world conversations  (Maloney et al., 2020; Sykownik et al., 2023; Wu et al., 2023), because virtual avatars reduce or even filter out several important non-verbal signals such as facial expressions and body language  (Baker et al., 2021; Freeman and Maloney, 2021; Zhang et al., 2022; Lee et al., 2022; Fu et al., 2023). As an extreme case, imperfect avatars with realism will lead to the uncanny valley effect, bringing discomfortable social experiences to users  (Latoschik et al., 2017; Kyrlitsias and Michael-Grigoriou, 2022). Besides the lack of expressiveness, the voice-based communication, as a common way to communicate in social VR, encounters the difficulty of coordinating speakers in multi-user virtual spaces  (Yan et al., 2023).
@@
However, people appreciate social VR not because it can simulate the real world. Instead, a significant reason is that the unique affordances of social VR can offer users immersive and unrealistic social experiences  (Freeman and Maloney, 2021; Maloney et al., 2020). In other words, social VR can augment social signaling and unlock new social interactions that are unattainable in the realistic world  (McVeigh-Schultz and Isbister, 2022, 2021). At present, the potential of virtual reality to create unrealistic experiences for enhancing interpersonal communication for socialization needs is still underexplored. This leaves us opportunities to dive deep into the potentials of virtual reality and build novel tools with the idea of “superpowers”  (McVeigh-Schultz and Isbister, 2022, 2021) to support interpersonal communication in social VR.
@@ 2.2. Combining Verbal and Non-verbal Information to Augment Communication
Both verbal and non-verbal information is crucial in interpersonal communication in social VR  (Palmer, 1995). Besides verbal cues that mainly convey semantic meanings, non-verbal cues convey emotions  (Liebman and Gergle, 2016; Luo et al., 2024), enhance understanding  (Aburumman et al., 2022), regulate interactions  (Maloney et al., 2020), reflect cultural norms  (Freeman and Acena, 2021), and overall enrich the communication experiences no matter in offline face-to-face scenarios or online digital spaces.
Recent research in the HCI and CSCW fields has explored ways to involve non-verbal signals into verbal-centric communications (e.g., live presentations) to facilitate interpersonal communication and build social connections through computational approaches  (Liao et al., 2022; Liu et al., 2023; Cao et al., 2024; An et al., 2024; Chen et al., 2021; Aoki et al., 2022; Choi and Aizawa, 2019). A thread of research aims to provide visual aids for computer-mediated spoken presentations. “RealityTalk”  (Liao et al., 2022) explores a speech-driven approach that allows users to give a speech with predefined visual elements while mentioning particular key words. “Visual Captions”  (Liu et al., 2023) and “CrossTalk”  (Xia et al., 2023) further automate the process by introducing machine-learning-based models to predict users’ intentions and automatically provide potential visual elements while processing speech input. “Elastica”  (Cao et al., 2024) further enables adaptive animations on the associated visuals to effectively enhance the expressiveness of presentations.
Beyond visual aids, affective enhancement is another important dimension explored by previous research. For voice messaging interfaces, “Emowear”  (An et al., 2024) invents a concept named “Emotional Teasers”, which refers to a collection of animated emoticons, to explicitly show the emotional tone of an upcoming voice message for smartwatch users. For text-centered communication mediums such as text messaging interfaces, previous research has explored how the color and shape of speech bubbles  (Chen et al., 2021; Aoki et al., 2022), and the typeface of textual content  (Choi and Aizawa, 2019; de Lacerda Pataca et al., 2023) can be purposefully designed and computationally generated to convey the speaker’s moods in conversations.
@@ 2.3. Impact Captions: A Typographic-Centered Visual Design Beyond Pure Verbal Communication
Impact captions refer to a type of typographic-centered visual effect that is prevalent in TV shows and online videos to engage viewers  (Sasamoto, 2014). Traditionally, impact captions are used as a supplementary channel that provides information outside of the screen (e.g., a commentary message from the TV show’s editor  (O’Hagan, 2010)) or visualizes the implicit non-verbal information in frames (e.g., the characters’ moods  (O’Hagan, 2010) or environmental sound effects  (Wang et al., 2016)). Unlike visual cues in previous related works for improving interpersonal communication  (Liao et al., 2022; Liu et al., 2023; Cao et al., 2024; An et al., 2024; Chen et al., 2021; Aoki et al., 2022; Choi and Aizawa, 2019), impact captions always use text as the main visual component that conveys verbal information and apply artistic modifications with additional visual decorations to textual elements to convey non-verbal cues simultaneously. The integration of both verbal and non-verbal information in the visual design of impact captions makes the captions a powerful medium that can provide rich semantics, afford viewers’ affective reactions, and provoke thoughts  (O’Hagan, 2010; Sasamoto et al., 2021; Chow, 2023).
@@ 3. Design Space of Impact Captions
As a form of typographic design as well as visual rhetoric, impact captions can convey both verbal and non-verbal information on multiple dimensions. To introduce impact captions into social VR as a communication medium, we investigated the design space of impact captions with a concentration on the visual components and interactions.
@@ 3.1. Methods: Analyzing Videos with Impact Captions
To obtain a comprehensive understanding of impact captions currently used in videos (e.g., TV shows and online videos), we conducted a content analysis on a collection of videos of TV show recordings retrieved from popular online video-sharing platforms.
@@ 3.1.1. Data Collection
We collected relevant videos using a two-stage approach. In the first stage, we followed the top TV series rankings (i.e., IMDb [^1] https://www.imdb.com/?ref_=nv_home, TV Time [^2] https://www.tvtime.com/, and Douban [^3] https://m.douban.com/tv/tvshow) to collect and review videos of three representative variety show series: “Arashi” (from Japan), “Running Man” (from South Korea), and “Who is the Murderer” (from China). These videos helped us get an initial sense and identify preliminary dimensions of the design space. In the second stage, we iteratively enhanced the design space by analyzing videos from an enlarged sample video set collected from online video-sharing platforms including YouTube and Bilibili. Specifically, we used the keywords “variety show with captions” and “popular variety shows” to search on these platforms with a newly registered account on YouTube and Bilibili, respectively. For each keyword, we focused on the top 30 videos of the raw results and filtered out irrelevant or redundant videos. As a result, a total of 46 sample videos were included in the dataset.

### v3

@@ Abstract.
Social Virtual Reality (VR) emerges as a promising platform which affords immersive, interactive, and engaging mechanisms for collaborative activities in virtual spaces. However, interpersonal communication in social VR is still limited with existing mediums and channels. To bridge the gap, we propose a novel method for mediating real-time conversations in social VR, which leverages impact captions, a type of typographic visual effect widely used in videos, to encode both verbal and non-verbal information. We first investigated the design space of impact captions by content analysis and a co-design session with four experts. We then implemented SpeechCap, a proof-of-concept system with which users can communicate with each other using speech-driven impact captions in VR. Through a user study (N=14), we evaluated the effectiveness of the visual and interaction design of speech-driven impact captions, highlighting their strengths in the interactivity and integrating verbal and non-verbal information in communication mediums. Finally, we discussed our main findings regarding visual rhetoric, interactivity, and ambiguity, and further provided design implications for facilitating interpersonal communication in social VR.
@@
copyright: ccjournalyear: 2025isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Virtual realityccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in interaction design
Figure 1. SpeechCap showcases a novel approach to mediate voice conversations in social VR by converting speech into playful impact captions. (A) From real-time speech, SpeechCap generates impact captions with customized visual design that conveys both verbal and non-verbal information. The caption “happy” is in a warm and bright color for the positive emotions, and a “smiling” emoji further highlight the pleasant and relaxing atmosphere. (B) Users can play with impact captions like real-world objects, e.g., throwing a “cake”. (C) Once the flying “cake” crashes on another user, it will explode like a firework.
@@ 1. Introduction
With the increasing availability of consumer-level virtual reality head-mounted displays (HMDs)  (Anthes et al., 2016; Li et al., 2021), Social Virtual Reality (VR) emerges as a promising platform that brings immersive, interactive, and engaging experiences to users for a wide range of collaborative activities. Social VR now supports remote conferencing  (Abdullah et al., 2021), immersive learning  (Thanyadit et al., 2022; Peng et al., 2022), healthcare  (Li et al., 2020; Udapola, 2022), collaborative design  (Mei et al., 2021), music composing  (Men and Bryan-Kinns, 2022) and even daily activities such as sleeping  (Maloney and Freeman, 2020) and drinking  (Chen et al., 2024). Among these diverse application areas, social VR helps users overcome physical dispersion, enabling them to communicate and collaborate in the same virtual environment  (Li et al., 2021; Palmer, 1995) by offering channels for both verbal and non-verbal communication  (Dzardanova et al., 2022; McVeigh-Schultz and Isbister, 2021; Wei et al., 2024). Specifically, text messaging and voice chat are two channels commonly used for verbal communication. And pictures  (Li et al., 2019), avatars  (Fu et al., 2023; Baker et al., 2021; Kolesnichenko et al., 2019; Freeman and Maloney, 2021), and embodied interactions  (Maloney et al., 2020; Wieland et al., 2022) are channels that can convey non-verbal information in social VR.
Yet, these existing communication channels in social VR are preliminary, as they are either borrowed from traditional media (e.g., texts, voices, and pictures) lacking adaptations to the nature of social VR  (Montoya Esquer and Lara López, 2023; Rzayev et al., 2021; Dzardanova et al., 2022; Hsieh et al., 2020), or difficult to be reproduced very well by current VR technologies (e.g., gestures, facial expressions, embodied actions)  (Sykownik et al., 2023; Tanenbaum et al., 2020; Wei et al., 2024). Specifically, texts as a communication medium in VR still suffer from difficulties in user inputting  (Montoya Esquer and Lara López, 2023) and readability issues in presentations  (Rzayev et al., 2021; Hsieh et al., 2020). Voice-based conversations in VR can easily be disturbed by sound intrusions from the environment  (Akselrad et al., 2023). Furthermore, non-verbal communication mediums, such as gestures and facial expressions, cannot be well-simulated with the avatars in current VR applications  (Baker et al., 2021; Aseeri and Interrante, 2021; Freeman and Maloney, 2021), and users may find it hard to achieve specific actions or movements using controllers with joysticks and buttons  (Sykownik et al., 2023; Tanenbaum et al., 2020; Li et al., 2019).
@@
To implement and evaluate our impact-caption-inspired approach, we first explored the visual and interaction design space of impact captions through a content analysis on TV show videos. Based on the design space, we held a co-design session with 4 experts (i.e., HCI researchers) to further derive specific design goals for building a system to demonstrate our idea. Following the findings so far, we developed SpeechCap, a proof-of-concept system that takes real-time speech as input to generate interactive impact captions with customized visual appearances in a multi-user VR environment. Using SpeechCap, we conducted an in-lab study with 14 participants to assess impact captions as a medium for supporting interpersonal communication in social VR.
@@
In summary, this research contributes: (i) A design space of impact captions for supporting interpersonal communication in social VR; (ii) A proof-of-concept system, SpeechCap, that enables users to have real-time voice-based conversations with the support of interactive impact captions in social VR; (iii) An evaluation study of SpeechCap that demonstrates the effectiveness of our impact-caption-inspired approach for communication, highlighting the value of employing creative mediums for supporting interpersonal communication in social VR.
@@ 2.1. Interpersonal Communication in Social Virtual Reality
Social Virtual Reality (VR) refers to immersive virtual spaces where multiple users can interact synchronously through VR head-mounted displays (HMDs)  (McVeigh-Schultz et al., 2019; Freeman and Maloney, 2021). It provides high-fidelity virtual presence to facilitate various forms of interpersonal communication with both verbal and non-verbal cues over other computer-mediated communication channels  (Yassien et al., 2020; McVeigh-Schultz and Isbister, 2021).
In an immersive and synchronous environment, social VR can effectively support collaborative tasks and promote social interactions in a wide range of application scenarios, including collaborative prototyping  (Mei et al., 2021), healthcare and treatment  (Li et al., 2020; Udapola, 2022), immersive learning  (Thanyadit et al., 2022; Peng et al., 2022; Jensen and Konradsen, 2018), intimate relationship building  (Wang et al., 2023; Freeman and Acena, 2021), and inter-generational communication  (Shen et al., 2024; Du et al., 2024; Wei et al., 2023; Baker et al., 2019).
Although social VR has been increasingly applied in various domains, it is still far from fulfilling users’ social and emotional needs as a digital space  (Wei et al., 2024; Tanenbaum et al., 2020; Maloney et al., 2020). Previous research identified several limitations regarding interpersonal communication in social VR. A known issue is that current social VR applications do not allow users to easily express emotions (e.g., mood and excitement) and intentions through non-verbal cues, which are common in real-world conversations  (Maloney et al., 2020; Sykownik et al., 2023; Wu et al., 2023), because virtual avatars reduce or even filter out several important non-verbal signals such as facial expressions and body language  (Baker et al., 2021; Freeman and Maloney, 2021; Zhang et al., 2022; Lee et al., 2022; Fu et al., 2023). As an extreme case, imperfect avatars with realism will lead to the uncanny valley effect, bringing discomfortable social experiences to users  (Latoschik et al., 2017; Kyrlitsias and Michael-Grigoriou, 2022). Besides the lack of expressiveness, voice-based communication, a common method in social VR, faces challenges in coordinating speakers in multi-user virtual spaces  (Yan et al., 2023).
@@
However, people appreciate social VR not because it can simulate the real world. Instead, a significant reason is that the unique affordances of social VR can offer users immersive and unrealistic social experiences  (Freeman and Maloney, 2021; Maloney et al., 2020). In other words, social VR can augment social signaling and unlock new social interactions that are unattainable in the realistic world  (McVeigh-Schultz and Isbister, 2022, 2021). At present, the potential of virtual reality to create unrealistic experiences for enhancing interpersonal communication for socialization needs is still under-explored. This leaves us opportunities to dive deep into the potentials of virtual reality and build novel tools with the idea of “superpowers”  (McVeigh-Schultz and Isbister, 2022, 2021) to support interpersonal communication in social VR.
@@ 2.2. Combining Verbal and Non-verbal Information to Augment Communication
Both verbal and non-verbal information is crucial in interpersonal communication in social VR  (Palmer, 1995). Besides verbal cues that mainly convey semantic meanings, non-verbal cues convey emotions  (Liebman and Gergle, 2016; Luo et al., 2024), enhance understanding  (Aburumman et al., 2022), regulate interactions  (Maloney et al., 2020), reflect cultural norms  (Freeman and Acena, 2021), and overall enrich communication experiences, whether in offline face-to-face scenarios or online digital spaces.
Recent research in the HCI and CSCW fields has explored ways to involve non-verbal signals in verbal-centric communications (e.g., live presentations) to facilitate interpersonal communication and build social connections through computational approaches  (Liao et al., 2022; Liu et al., 2023; Cao et al., 2024; An et al., 2024; Chen et al., 2021; Aoki et al., 2022; Choi and Aizawa, 2019). A thread of research aims to provide visual aids for computer-mediated spoken presentations. “RealityTalk”  (Liao et al., 2022) explores a speech-driven approach that allows users to give a speech with predefined visual elements while mentioning particular key words. “Visual Captions”  (Liu et al., 2023) and “CrossTalk”  (Xia et al., 2023) further automate the process by introducing machine-learning-based models to predict users’ intentions and automatically provide potential visual elements while processing speech input. “Elastica”  (Cao et al., 2024) further enables adaptive animations on the associated visuals to effectively enhance the expressiveness of presentations.
Beyond visual aids, enhancing emotional communication is another important dimension explored by previous research. For voice messaging interfaces, “Emowear”  (An et al., 2024) introduced the concept of “Emotional Teasers,” a collection of animated emoticons designed to show the emotional tone of upcoming voice messages for smartwatch users. For text-centered communication mediums such as text messaging interfaces, previous research has explored how the color and shape of speech bubbles  (Chen et al., 2021; Aoki et al., 2022), and the typeface of textual content  (Choi and Aizawa, 2019; de Lacerda Pataca et al., 2023) can be purposefully designed and computationally generated to convey the speaker’s moods in conversations.
@@ 2.3. Impact Captions: A Typographic-Centered Visual Design Beyond Pure Verbal Communication
Impact captions refer to a type of typographic-centered visual effect that is prevalent in TV shows and online videos to engage viewers  (Sasamoto, 2014). Traditionally, impact captions are used as a supplementary channel to provide information outside the screen (e.g., commentary messages from the TV show’s editor  (O’Hagan, 2010)) or to visualize implicit non-verbal information within frames (e.g., characters’ moods or environmental sound effects  (Wang et al., 2016)). Unlike visual cues in previous work for improving interpersonal communication  (Liao et al., 2022; Liu et al., 2023; Cao et al., 2024; An et al., 2024; Chen et al., 2021; Aoki et al., 2022; Choi and Aizawa, 2019), impact captions always use text as the main visual component that conveys verbal information and apply artistic modifications with additional visual decorations to textual elements to convey non-verbal cues simultaneously. The integration of both verbal and non-verbal information in the visual design of impact captions makes the captions a powerful medium that can provide rich semantics, evoke emotional reactions in viewers, and provoke thoughts  (O’Hagan, 2010; Sasamoto et al., 2021; Chow, 2023).
@@ 3. Design Space of Impact Captions
In order to introduce impact captions into social VR as a medium for interpersonal communication, we explored the design space of impact captions through a content analysis on the TV variety show videos. Our design space consists of the visual design (Figure 2) and the interaction design (Figure 3) of impact captions. Furthermore, the visual design space separately considered textual elements and non-textual elements.
@@ 3.1. Methods: Analyzing Videos with Impact Captions
Three researchers conducted the content analysis collaboratively on a collection of videos of TV variety show, in which impact captions are commonly used, to explore the design of impact captions.
@@ 3.1.1. Data Collection
We collected relevant videos using a two-stage approach. Firstly, we followed the top TV series rankings (i.e., IMDb [^1] https://www.imdb.com/?ref_=nv_home, TV Time [^2] https://www.tvtime.com/, and Douban [^3] https://m.douban.com/tv/tvshow) to select TV shows where impact captions are widely used. Specifically, we concentrated on three representative variety show series: “Arashi” (from Japan), “Running Man” (from South Korea), and “Who is the Murderer” (from China). For each show, we reviewed one randomly-selected episode. Secondly, we enlarged the collection by including videos found from YouTube and Bilibili, which are two popular online video-sharing platforms. On both platforms, we used the keywords “variety show with captions” and “popular variety shows” (we translated the keywords into Chinese version on Bilibili) to search for videos using a newly registered account. From the raw results of each keyword, we took the top-30 videos and filtered out irrelevant or redundant videos. Finally, 46 sample videos were included.